OpenStudy (anonymous):

l = 8, w = 4, h = 2 Find the diagonal (d) of the base.

OpenStudy (anonymous):



OpenStudy (anonymous):

What's the formula?

OpenStudy (anonymous):

it gives me ___=____√____________ thats the answer format

OpenStudy (anonymous):

@CrazyCountryGirl

OpenStudy (anonymous):

...I don't know how to do this, I'm so so sorry!......

OpenStudy (anonymous):

It's basic trig if that helps

OpenStudy (anonymous):

Wont the pythagorean theorum solve this? 8^2+4^2=x^2

OpenStudy (anonymous):

Im not sure which base they want to you solve for..

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

are you sure that is trignometry? you are just finding the diagonal of the base which is L x W you can just use the pythagorean theorum to find the diagonal

OpenStudy (acxbox22):



OpenStudy (anonymous):

we dont have to use trig, but it is the quiz for the trig section so I just thought it might be

OpenStudy (acxbox22):

so.... a^2+b^2=c^2 8^2+4^2=c^2 64+16=c^2 80=c^2 \[c=\sqrt{80}=\sqrt{16\times5}=4\sqrt{5}\]

OpenStudy (anonymous):

thank you, idk why but I always over think things and mess up :p